OpenStudy (anonymous):

solve; 2x^2+12x+16=0

OpenStudy (lgbasallote):

\[2x^2 + 12x + 16 = 0\] \[2(x^2 + 6x + 8) = 0\] \[2(x+4)(x+2) = 0\] therefore \[x = -4\] \[x = -2\]